# QueStER: Query Specification for Generative Retrieval

### v1

@@ Abstract
Generative Retrieval (GR) differs from the traditional index–then–retrieve pipeline by storing relevance in model parameters and directly generating document identifiers. However, GR often struggles to generalize and is costly to scale. We introduce QueStER (QUEry SpecificaTion gEnerative Retrieval), which reframes GR as query specification generation – in this work, a simple keyword query handled by BM25 – using a (small) LLM. The policy is trained using reinforcement learning techniques (GRPO). Across in- and out-of-domain evaluations, we show that our model is more effective than BM25, and competitive with neural IR models, while maintaining a good efficiency.
QueStER: Query Specification for Generative Retrieval
Arthur Satouf1,2,3,4,5          Yuxuan ZONG5
Habiboulaye Amadou-Boubacar3          Pablo Piantanida1,4          Benjamin Piwowarski5
Preprint — under review
1 MILA – Quebec AI Institute, ILLS, Canada
2 Université Paris-Saclay, France
3 Air Liquide, France
4 CentraleSupélec, CNRS, France
5 Sorbonne Université, ISIR, CNRS, France
arthur.satouf@gmail.comzong@isir.upmc.frbenjamin.piwowarski@cnrs.frhabiboulaye.amadou-boubacar@airliquide.compablo.piantanida@cnrs.fr
@@ 1 Introduction
Information Retrieval (IR) models aim to retrieve relevant documents from a large database that satisfy a user information need. They are a key component of search engines and of Retrieval-Augmented Generation (RAG) models (lewis2020retrieval). For a long time, IR BoW (bag-of-words) models, such as BM25 robertson1994okapi, have been based on term-specific statistics, allowing fast retrieval through the use of efficient index structures such as Block-Max WAND grand_maxscore_2020. However, bag-of-words models suffer from the vocabulary mismatch problem, whereby the user query might contain keywords that do not appear in relevant documents.
@@
A standard way to solve this problem is to rely on query rewriting techniques. The first works lavrenko2001relevance; abdul-jaleel2004umass add keywords extracted from the documents retrieved by the original query. This might lead to query drift, especially if the retrieved documents are not related to the user information need. More recently, jagerman2023query; alaofi2023can leveraged LLMs to improve the quality of the rewritten query. While improving compared to previous works, these strategies degrade when using a small LLM (e.g. smaller than 4B of parameters). When using large LLMs (with complex prompts), and moreover sampling several times (zhang2024exploring), which decrease the efficiency. Exploring alternatives that balance effectiveness and efficiency is important.
@@
Nowadays, a common alternative to increase the effectiveness of IR models is to build models on top of Transformer (vaswani2017transformer) architectures. Among the different models, dual encoders (dense and sparse), as well as late-interaction models such as ColBERT (santhanam2021colbertv2), are the most effective and efficient first stage neural rankers. Other architectures such as cross-encoders (nogueira2019passage) are only used for re-ranking, and are not the focus of this paper. However, these neural IR models have shortcomings since using a neural (first-stage) IR model efficiently requires indices (either dense or sparse, depending on the type of model). This is costly for two reasons. First, the space taken by those indices is much larger than that for BoW models (e.g. a dense index on MS MARCO with 8.8M documents using FP16 for 768-dimensional vectors requires 13 GiB, while the BM25 index only takes 0.67 GiB (park2025decoding)). Second, when the model is re-trained, the index must be rebuilt, which is not practical with large collections such as that of the OpenSearch Euro project (granitzer2024impact).
@@
To tackle this issue, Generative Retrieval (GR) models attempt to internalize the index in the parameters (tay2022transformer), which avoids entirely the construction of an index. During inference, these models directly predict a document identifier conditioned on the user query. Some GR models rely on arbitrary document identifiers (tay2022transformer; wang2022neural), but these models do not generalize well, especially for large collections. Another research direction is to use document meta-data (zhou2022ultron; tang2023semantic; zhang2024generative) – such as URL, title, or some keywords from the document – as identifiers. While these approaches generalize better, they are still limited by the metadata they use.
@@
In this work, we propose a model named QueStER (for Search Specification Generative Retrieval). QueStER is based on an LLM whose parameters are optimized, with GRPO (shao2024deepseekmath), to generate keyword queries for a search engine. In this work, we restrict the search engine to be a keyword-based BM25. Through our experiments, we show that QueStER is both effective (+4.0 nDCG@10 in-domain and +5.3 out-of-domain, compared to BM25) and efficient ($\approx 28$ ms/query for BM25). QueStER reaches the performance of prompt-based methods that depend on much larger LLMs, despite using only a 4B backbone LLM.
@@ 2 Related Work
Information Retrieval focuses on efficiently and effectively retrieving relevant documents that satisfy users information needs. Although current neural approaches such as DPR (karpukhin2020dense) or RepLLaMA (ma2023finetuning) dominate the leaderboards, traditional sparse models such as BM25 (robertson1994okapi) are still widely considered by the community due to their excellent tradeoff between efficiency and effectiveness.
@@
Models retrieve documents by directly generating their identifiers (DocIDs) using language models, bypassing the indexing stage. A first series of work  (tay2022transformer; wang2022neural; sun2023learning; yang2023auto; lee2023glen) proposes to associate with a document an arbitrary DocID which comes from a (hierarchical) clustering process or is learned. In both cases, these models have limited power of generalization and their effectiveness decreases with the number of documents in the collection. An alternative is to use “natural” DocIDs, such as document titles (de2020autoregressive; chen2022gere; chen2022corpusbrain), URLs (zhou2022ultron), N-grams (bevilacqua2022autoregressive) or document summaries (tang2023semantic). Although this offers good interpretability, meta-data are not a perfect representation of document semantic content, and the performance of such models still lags compared to other neural approaches on large-scale datasets (pradeep2023does).
@@
Query rewriting (carpineto2012survey) is a common solution in IR to improve query precision and expressiveness. Statistical models such as BM25 robertson1994okapi use pseudo-relevance feedback methods (lavrenko2001relevance; abdul-jaleel2004umass) that rely on heuristics and/or statistical analysis of the initial query with its retrieved documents. These methods are often prone to query drift (mitra1998improving), which limits their applicability. Taking advantage of the performance of recent LLMs (yang2025qwen3; grattafiori2024llama; team2025gemma), various LLM-based query rewriting methods have been proposed. Many works directly leverage the strong capabilities of LLMs to expand or rewrite queries through various prompting strategies, including natural language questions (alaofi2023can; ye2023enhancing), keywords (jagerman2023query; li2024query; mackie2023generative) or even passages (gao2023precise; wang2023query2doc; shen2024large; zhang2024exploring). While this leads to an increased retrieval effectiveness, successful models rely on complex prompts, very large models, and multiple samples, each being costly at inference time. These drawbacks led to the development of approaches more closely related to ours, and which propose to finetune smaller LLMs. Following (nogueira2017task), several works (mao2024rafe; ma2023query; peng2024large; hsu2024grounding; yao2025llm) propose to leverage reinforcement learning techniques using a multi-source reward – e.g. annotated samples, cross-encoder like MonoBERT (nogueira2019passage). However, these methods are used within a multi-hop scenario (hsu2024grounding), with low efficiency, or are applied in a RAG system (mao2024rafe), and did not really look the underlying IR performance (as well as using a weaker reward signal as in our work).
@@ 3.1 Problem Formulation
Our goal is to build a model able to generate a query specification $p_{\theta}(q)$ from an initial user query $q$, where the re-written query leads to both efficient and effective retrieval. To process the query specification $p_{\theta}(q)$, we use existing BoW search engines $se$, e.g. BM25 robertson1994okapi, to process the search specification $p_{\theta}(q)$ depending on the search engine $se$. In this work, we use a bag-of-words model, namely BM25.
@@ 3.2 Policy Optimization
We optimize the rewriting policy $p_{\theta}$ with Group-Relative Policy Optimization (GRPO(shao2024deepseekmath). For each query $q$, the policy samples a group of $m$ candidates $\{q_{i}\}_{i=1}^{m}$, their associated rewards $\{r_{i}\}$ are calculated as described in Section 3.3, and group-relative advantages $a_{i}=r_{i}-\bar{r}$ (with $\bar{r}=\tfrac{1}{m}\sum_{i}r_{i}$) drive a clipped policy-gradient update. Following prior large-scale GRPO practice (hu2025open), we standardize rewards within each group and set the KL weight to $\beta=0$ (omit an explicit KL penalty) to encourage exploration (ablations in Section 4 justify this choice).
@@ 3.3.1 Soft and Hard nDCG
An IR metric that matches our needs is SoftRank, which computes the expected nDCG taylor2008softrank as
@@
An alternative metric would be to use the InfoNCE loss oord2018representation, commonly used to train neural IR models. However, in our preliminary experiments the training was unstable compared to the one using SoftRank (taylor2008softrank).
@@ 3.3.2 Distillation: Cross-Encoder Assessments
The dataset used most often to train IR models is MS MARCO (bajaj2016ms), where the evaluations of whether a document is relevant to a given user query are based on user clicks. User clicks are known to be incomplete Gupta_2022: many queries are associated with a single relevant document, i.e. MS MARCO contains many false negatives. In our case, $\mathbb{E}(nDCG)$ can be zero for all query candidates, which cannot be used for GRPO to learn properly. To tackle this issue, we use distillation hofstatter_efficiently_2021, a standard technique used to train IR models that takes advantage of the most powerful IR models, namely cross-encoders (CE), to indicate whether or not a document is relevant.
@@ 4.1 Experimental Setup
We train our model on the MS MARCO v1 passage dataset (bajaj2016ms). This dataset contains 8.8M passages. We evaluate the model in-domain on the MS MARCO dev set (6980 queries), as well as the TREC DL 2019 (43 assessed topics) and TREC DL 2020 (54 assessed topics) sets. Following common practice, we report out-of-domain (OOD) performance of our method using the BEIR dataset (thakur2021beir).
@@
  • ANCE (karpukhin2020dense), a state-of-the-art dense neural IR model;
  • SPLADEv2 (formal2021splade), a state-of-the-art sparse neural IR model;
  • ColBERTv2 (santhanam2021colbertv2), a state-of-the-art multi-vector retrieval model, which is especially good at generalization.
@@
  • A standard IR baseline, BM25 robertson1994okapi, which also serve as the retrieving model for our rewritten queries.
  • RM3 lavrenko2001relevance, a query rewriting model based on relevance feedback for BM25 retrieval.
  • HyDE gao2023precise, a query rewriting model based on the generation of relevant passages;
  • Query2Doc wang2023query2doc (abbr. Q2D), similar to HyDE, but that concatenates the query for BM25 retrieval.
  • LameR shen2024large, a LLM-based query rewriting model that concatenates the top retrieved documents.
  • MuGI zhang2024exploring, a state-of-the-art work that balances the query and pseudo-document for query rewriting. We report both the result based on ChatGPT3.5 and ChatGPT4.
The results for SPLADEv2 and ColBERTv2 are copied from  (santhanam2021colbertv2). The results for ANCE, RM3 and Q2D are copied from (wang2023query2doc). The other baselines are from (zhang2024exploring). We report very few generative retrieval baseline results as they do not usually evaluate on BEIR and on the full MS MARCO dataset. [^2] We only report MINDER (li2023multiview) for which a few results on full MS MARCO are reported. Among all the results reported in these papers, we generally select the ones reported in the tables by using the configuration put forward by the author. For the neural dual encoders, we take the result of the model trained with distillation (hofstatter_efficiently_2021), as we are also distilling from a cross-encoder.
@@
We conduct all our experiments [^3] Huggingface Transformers 4.53.3, TRL 0.19.1 with NVIDIA RTX A6000 48GB gpus (total cost \$150 - \$200). For GRPO training, we fine-tune Qwen3 models yang2025qwen3 model [^4] QWEN3 across various size (0.6B, 1.7B and 4B) with LoRA (hu2022lora) (rank $r=40$, scaling $\alpha=40$, “thinking” mode disabled). The 4B version is our main model as it offers the best efficiency/effectiveness trade-off. For each input query, we sample from the LLM a group of $10$ candidates with a temperature $\tau=1.2$. The model is trained on 96,00096{,}000 randomly sampled queries from MS MARCO training set. As for the assessments, we either use the original assessments from MS MARCO train set, or the cross-encoder ones from OpenSearch  [^5] OpenSearch Project – MS MARCO Hard Negatives LLM Scores. Cross-Encoder-based relevance assessments are normalized by dividing the maximum score by the maximum score for a given query, leading to values in the range $[0,1]$.

### v2

# QueStER: Query Specification for Generative Keyword-Based Retrieval

@@ Abstract
Generative retrieval (GR) differs from the traditional index–then–retrieve pipeline by storing relevance in model parameters and generating retrieval cues directly from the query, but it can be brittle out of domain and expensive to scale. We introduce QueStER (QUEry SpecificaTion gEnerative Keyword-Based Retrieval), which bridges GR and query reformulation by learning to generate explicit keyword-based search specifications. Given a user query, a lightweight LLM produces a keyword query that is executed by a standard retriever (BM25), combining the generalization benefits of generative query rewriting with the efficiency and scalability of lexical indexing. We train the rewriting policy with reinforcement learning techniques. Across in- and out-of-domain evaluations, QueStER consistently improves over BM25 and is competitive with neural IR baselines, while maintaining strong efficiency. [^1] Code on Github.
QueStER: Query Specification for Generative Keyword-Based Retrieval
Arthur Satouf1,2,3,4, Yuxuan Zong4, Habiboulaye Amadou-Boubacar3, Pablo Piantanida1,2, Benjamin Piwowarski4 1 MILA – Quebec AI Institute & ILLS, Canada 2Université Paris-Saclay & CentraleSupélec & CNRS, France 3Air Liquide, France 4Sorbonne Université & ISIR & CNRS, France arthur.satouf(at)gmail.com, zong@isir.upmc.fr, habiboulaye.amadou-boubacar@airliquide.com, pablo.piantanida@cnrs.fr, benjamin.piwowarski@cnrs.fr
@@ 1 Introduction
Information Retrieval (IR) models aim to retrieve relevant documents from a large database that satisfy a user information need. They are a key component of search engines and of Retrieval-Augmented Generation (RAG) models (Lewis et al., 2020). For a long time, IR BoW (bag-of-words) models, such as BM25 Robertson et al. (1994), have been based on term-specific statistics, allowing fast retrieval through the use of efficient index structures such as Block-Max WAND Grand et al. (2020). However, bag-of-words models suffer from the vocabulary mismatch problem, whereby the user query might contain keywords that do not appear in relevant documents.
@@
A standard way to solve this problem is to rely on query rewriting techniques. The first works Lavrenko and Croft (2001); Abdul-Jaleel et al. (2004) add keywords extracted from the documents retrieved by the original query. This might lead to query drift, especially if the retrieved documents are not related to the user information need. More recently, Jagerman et al. (2023); Alaofi et al. (2023) leveraged LLMs to improve the quality of the rewritten query. While improving compared to previous works, these strategies degrade when using a small LLM (e.g. smaller than 4B of parameters). When using large LLMs (with complex prompts), and moreover sampling several times (Zhang et al., 2024a), which decrease the efficiency. Exploring alternatives that balance effectiveness and efficiency is important.
Nowadays, a common alternative to increase the effectiveness of IR models is to build models on top of Transformer (Vaswani et al., 2017) architectures. Among the different models, dual encoders (dense and sparse), as well as late-interaction models such as ColBERT (Santhanam et al., 2021), are the most effective and efficient first stage neural rankers. Other architectures such as cross-encoders (Nogueira and Cho, 2019) are only used for re-ranking, and are not the focus of this paper. In addition, learned sparse lexical retrievers (e.g., DeepImpact, RRA) (Basnet et al., 2024; Satouf et al., 2025) predict context-dependent term weights and sometimes lexical expansion, mitigating vocabulary mismatch while retaining inverted-index retrieval.
@@
However, neural IR models have shortcomings since using a neural (first-stage) IR model efficiently requires indices (either dense or sparse, depending on the type of model). This is costly for two reasons. First, the space taken by those indices is much larger than that for BoW models (e.g. a dense index on MS MARCO with 8.8M documents using FP16 for 768-dimensional vectors requires 13 GiB, while the BM25 index only takes 0.67 GiB (Park et al., 2025)). Second, when the model is re-trained, the index must be rebuilt, which is not practical with large collections such as that of the OpenSearch Euro project (Granitzer et al., 2024).
@@
To tackle this issue, Generative Retrieval (GR) models attempt to internalize the index in the parameters (Tay et al., 2022), which avoids entirely the construction of an index. During inference, these models directly predict a document identifier conditioned on the user query. Some GR models rely on arbitrary document identifiers (Tay et al., 2022; Wang et al., 2022), but these models do not generalize well, especially for large collections. Another research direction is to use document meta-data (Zhou et al., 2022; Tang et al., 2023; Zhang et al., 2024b) – such as URL, title, or some keywords from the document – as identifiers. While these approaches generalize better, they are still limited by the metadata they use.
@@
In this work, we propose a model named QueStER (for Search Specification Generative Retrieval). QueStER is based on an LLM whose parameters are optimized, with GRPO (Shao et al., 2024), to generate keyword queries for a search engine. In this work, we restrict the search engine to be a keyword-based BM25. Through our experiments, we show that QueStER is both effective (+4.0 nDCG@10 in-domain and +5.3 out-of-domain, compared to BM25) and efficient ($\approx 28$ ms/query for BM25). QueStER reaches the performance of prompt-based methods that depend on much larger LLMs, despite using only a 4B backbone LLM.
@@ 2 Related Work
Information Retrieval focuses on efficiently and effectively retrieving relevant documents that satisfy users information needs. Although current neural approaches such as DPR (Karpukhin et al., 2020) or RepLLaMA (Ma et al., 2024) dominate the leaderboards, traditional sparse models such as BM25 (Robertson et al., 1994) are still widely considered by the community due to their excellent tradeoff between efficiency and effectiveness.
@@
Models retrieve documents by directly generating their identifiers (DocIDs) using language models, bypassing the indexing stage. A first series of work  (Tay et al., 2022; Wang et al., 2022; Sun et al., 2023; Yang et al., 2023; Lee et al., 2023) proposes to associate with a document an arbitrary DocID which comes from a (hierarchical) clustering process or is learned. In both cases, these models have limited power of generalization and their effectiveness decreases with the number of documents in the collection. An alternative is to use “natural” DocIDs, such as document titles (De Cao et al., 2020; Chen et al., 2022a, b), URLs (Zhou et al., 2022), N-grams (Bevilacqua et al., 2022) or document summaries (Tang et al., 2023). Although this offers good interpretability, meta-data are not a perfect representation of document semantic content, and the performance of such models still lags compared to other neural approaches on large-scale datasets (Pradeep et al., 2023).
@@
Query rewriting (Carpineto and Romano, 2012) is a common solution in IR to improve query precision and expressiveness. Statistical models such as BM25 Robertson et al. (1994) use pseudo-relevance feedback methods (Lavrenko and Croft, 2001; Abdul-Jaleel et al., 2004) that rely on heuristics and/or statistical analysis of the initial query with its retrieved documents. These methods are often prone to query drift (Mitra et al., 1998), which limits their applicability. Taking advantage of the performance of recent LLMs (Yang et al., 2025; Grattafiori et al., 2024; Team et al., 2025), various LLM-based query rewriting methods have been proposed. Many works directly leverage the strong capabilities of LLMs to expand or rewrite queries through various prompting strategies, including natural language questions (Alaofi et al., 2023; Ye et al., 2023), keywords (Jagerman et al., 2023; Li et al., 2024; Mackie et al., 2023) or even passages (Gao et al., 2023; Wang et al., 2023; Shen et al., 2024; Zhang et al., 2024a). While this leads to an increased retrieval effectiveness, successful models rely on complex prompts, very large models, and multiple samples, each being costly at inference time. These drawbacks led to the development of approaches more closely related to ours, and which propose to finetune smaller LLMs. Following (Nogueira and Cho, 2017), several works (Mao et al., 2024; Ma et al., 2023; Peng et al., 2024; Hsu et al., 2024; Yao et al., 2025) propose to leverage reinforcement learning techniques using a multi-source reward – e.g. annotated samples, cross-encoder like MonoBERT (Nogueira and Cho, 2019). However, these methods are used within a multi-hop scenario (Hsu et al., 2024), with low efficiency, or are applied in a RAG system (Mao et al., 2024), and did not really look the underlying IR performance (as well as using a weaker reward signal as in our work).
@@ 3.1 Problem Formulation
Our goal is to build a model able to generate a query specification $p_{\theta}(q)$ from an initial user query $q$, where the re-written query leads to both efficient and effective retrieval. To process the query specification $p_{\theta}(q)$, we use existing BoW search engines $se$, e.g. BM25 Robertson et al. (1994), to process the search specification $p_{\theta}(q)$ depending on the search engine $se$. In this work, we use a bag-of-words model, namely BM25.
@@ 3.2 Policy Optimization
We optimize the rewriting policy $p_{\theta}$ with Group-Relative Policy Optimization (GRPO(Shao et al., 2024). For each query $q$, the policy samples a group of $m$ candidates $\{q_{i}\}_{i=1}^{m}$, their associated rewards $\{r_{i}\}$ are calculated as described in Section 3.3, and group-relative advantages $a_{i}=r_{i}-\bar{r}$ (with $\bar{r}=\tfrac{1}{m}\sum_{i}r_{i}$) drive a clipped policy-gradient update. Following prior large-scale GRPO practice (Hu et al., 2025), we standardize rewards within each group and set the KL weight to $\beta=0$ (omit an explicit KL penalty) to encourage exploration (ablations in Section 4 justify this choice).
@@ 3.3.1 Soft and Hard nDCG
An IR metric that matches our needs is SoftRank, which computes the expected nDCG Taylor et al. (2008) as
@@
An alternative metric would be to use the InfoNCE loss Oord et al. (2018), commonly used to train neural IR models. However, in our preliminary experiments the training was unstable compared to the one using SoftRank (Taylor et al., 2008).
@@ 3.3.2 Distillation: Cross-Encoder Assessments
The dataset used most often to train IR models is MS MARCO (Bajaj et al., 2016), where the evaluations of whether a document is relevant to a given user query are based on user clicks. User clicks are known to be incomplete Gupta and MacAvaney (2022): many queries are associated with a single relevant document, i.e. MS MARCO contains many false negatives. In our case, $\mathbb{E}(nDCG)$ can be zero for all query candidates, which cannot be used for GRPO to learn properly. To tackle this issue, we use distillation Hofstätter et al. (2021), a standard technique used to train IR models that takes advantage of the most powerful IR models, namely cross-encoders (CE), to indicate whether or not a document is relevant.
@@ 4.1 Experimental Setup
We train our model on the MS MARCO v1 passage dataset (Bajaj et al., 2016). This dataset contains 8.8M passages. We evaluate the model in-domain on the MS MARCO dev set (6980 queries), as well as the TREC DL 2019 (43 assessed topics) and TREC DL 2020 (54 assessed topics) sets. Following common practice, we report out-of-domain (OOD) performance of our method using the BEIR dataset (Thakur et al., 2021).
@@
  • ANCE (Karpukhin et al., 2020), a state-of-the-art dense neural IR model;
  • SPLADEv2 (Formal et al., 2021), a state-of-the-art sparse neural IR model;
  • ColBERTv2 (Santhanam et al., 2021), a state-of-the-art multi-vector retrieval model, which is especially good at generalization.
@@
  • A standard IR baseline, BM25 Robertson et al. (1994), which also serve as the retrieving model for our rewritten queries.
  • RM3 Lavrenko and Croft (2001), a query rewriting model based on relevance feedback for BM25 retrieval.
  • HyDE Gao et al. (2023), a query rewriting model based on the generation of relevant passages;
  • Query2Doc Wang et al. (2023) (abbr. Q2D), similar to HyDE, but that concatenates the query for BM25 retrieval.
  • LameR Shen et al. (2024), a LLM-based query rewriting model that concatenates the top retrieved documents.
  • MuGI Zhang et al. (2024a), a state-of-the-art work that balances the query and pseudo-document for query rewriting. We report both the result based on ChatGPT3.5 and ChatGPT4.
The results for SPLADEv2 and ColBERTv2 are copied from  (Santhanam et al., 2021). The results for ANCE, RM3 and Q2D are copied from (Wang et al., 2023). The other baselines are from (Zhang et al., 2024a). We report very few generative retrieval baseline results as they do not usually evaluate on BEIR and on the full MS MARCO dataset. [^3] We only report MINDER (Li et al., 2023) for which a few results on full MS MARCO are reported. Among all the results reported in these papers, we generally select the ones reported in the tables by using the configuration put forward by the author. For the neural dual encoders, we take the result of the model trained with distillation (Hofstätter et al., 2021), as we are also distilling from a cross-encoder.
@@
We conduct all our experiments [^4] Huggingface Transformers 4.53.3, TRL 0.19.1 with NVIDIA RTX A6000 48GB gpus (total cost \$150 - \$200). For GRPO training, we fine-tune Qwen3 models Yang et al. (2025) model [^5] QWEN3 across various size (0.6B, 1.7B and 4B) with LoRA (Hu et al., 2022) (rank $r=40$, scaling $\alpha=40$, “thinking” mode disabled). The 4B version is our main model as it offers the best efficiency/effectiveness trade-off. For each input query, we sample from the LLM a group of $10$ candidates with a temperature $\tau=1.2$. The model is trained on 96,00096{,}000 randomly sampled queries from MS MARCO training set. As for the assessments, we either use the original assessments from MS MARCO train set, or the cross-encoder ones from OpenSearch  [^6] OpenSearch Project – MS MARCO Hard Negatives LLM Scores. Cross-Encoder-based relevance assessments are normalized by dividing the maximum score by the maximum score for a given query, leading to values in the range $[0,1]$.